OpenStudy (anonymous):

help!!! list all the sets that -3 3/4 belongs to. a) natural numbers b) whole numbers c) intergers d) rational numbers E)irrational numbers f) real numbers

jimthompson5910 (jim_thompson5910):

-3 3/4 is a rational number since -3 3/4 = -15/4 -3 3/4 is a real number So the answers are D and F

OpenStudy (anonymous):



OpenStudy (anonymous):

and which one does 8 belong to? (different question)

jimthompson5910 (jim_thompson5910):

8 is an integer

jimthompson5910 (jim_thompson5910):

8 is a natural number

jimthompson5910 (jim_thompson5910):

8 is a whole number

jimthompson5910 (jim_thompson5910):

8 is rational since 8 = 8/1 and 8 is a real number

OpenStudy (anonymous):

THANK YOU<3

jimthompson5910 (jim_thompson5910):

so it's everything but choice E)